OpenStudy (anonymous):

Find sin θ, sec θ, and tan θ for an angle in standard position having this point on its terminal side: (/[−3,2sqrt{2}])

OpenStudy (anonymous):

Find sin θ, sec θ, and tan θ for an angle in standard position having this point on its terminal side: \[(-3, 2\sqrt{2})\]

OpenStudy (mertsj):



OpenStudy (mertsj):

\[\sin \theta=\frac{2\sqrt{2}}{\sqrt{17}}\]

OpenStudy (mertsj):

\[\tan \theta=-\frac{2\sqrt{2}}{3}\]

OpenStudy (mertsj):

\[\sec \theta=-\frac{\sqrt{17}}{3}\]